Mrs. Huang operates a soybean fam. She buys many supplies in bulk. Often the bulk products need to be custom mixed before Mrs. Huang can use them. To apply herbicide to a large field she must mix a solution of 67% herbicide with a solution of 46% herbicide to form 42 liters of a 55% solution. How much of the 67% solution must she use?

a. 23L b. 18L c. 34L d. 35L

Answer by ptaylor(2198)

Let x=amount of 67% solution needed
Then 42-x=amount of 46% solution needed
Now we know that the pure herbicide in the 67% solution (0.67x) plus the amount of pure herbicide in the 46% solution (0.46(42-x)) has to equal the amount of pure herbicide in the final mixture (0.55*42). So, our equation to solve is:
0.67x+0.46(42-x)=0.55*42 simplify
0.67x+19.32-0.46x=23.1 subtract 19.32 from each side
0.67x+19-32-19.32-0.46x=23.1-19.32 collect like terms
0.21x=3.78 divide each side by 0.21
x=18 liters------amount of 67% solution needed
Hope this helps---ptaylor